OpenStudy (anonymous):

medal If f(x) = 2x - 4, and g(x) = (x^2 + x + 5), find g (f(x))

OpenStudy (babyslapmafro):

Replace any "x" in the function g(x) with f(x)

OpenStudy (babyslapmafro):

for example:

OpenStudy (babyslapmafro):

if f(x) = x+2 and g(x) = x^2+4 g(f(x))= (x+2)^2 + 4

OpenStudy (anonymous):

x) = x+2 and g(x) = x^2+4 g(f(x))= (x+2)^2 + 4 @Babyslapmafro this is the right equation.

OpenStudy (babyslapmafro):

@aballer101 is that clear?

OpenStudy (anonymous):

g (f(x)) = g(2x-4) (substituted f(x)=2x-4 ) g(x) = (x^2 + x + 5) g(2x-4) means instead of x we have to substitute 2x-4 in g(x) g(2x-4)=(2x-4)^2+(2x-4)+5

OpenStudy (babyslapmafro):

correct

OpenStudy (anonymous):

a=-2x and b=4

OpenStudy (anonymous):

g(2x-4)=(2x-4)^2+(2x-4)+5

OpenStudy (babyslapmafro):

now just simplify that equation

OpenStudy (anonymous):

thank you g=fpor the report

OpenStudy (anonymous):

Your Welcome :)

OpenStudy (anonymous):

u did it ? @TypnoEchosis

OpenStudy (anonymous):

Indeed I did.

OpenStudy (anonymous):

2gx-4g-4x2+14x-17 = 0 is that right?

OpenStudy (anonymous):

why is that? cant have some fun?

OpenStudy (babyslapmafro):

simplify this part of the equation: (2x-4)^2+(2x-4)+5

OpenStudy (anonymous):

4x2 - 14x + 17

OpenStudy (anonymous):

I am sorry and it wont happen again but there was no need to report me

OpenStudy (anonymous):

@grrrxxiii i didnt report u

OpenStudy (anonymous):

i know it was the @TypnoEchosis

OpenStudy (babyslapmafro):

expand (2x-4)^2 first: (2x-4)^2 = 4x^2 - 16x + 16

OpenStudy (babyslapmafro):

yes, you have the correct final answer: 4x2 - 14x + 17

OpenStudy (anonymous):

thats it.. i dont have to simplify 4x2 - 14x + 17 ?

OpenStudy (babyslapmafro):

i would leave it in this form. a quadratic equation.

OpenStudy (anonymous):

aight thanks fam

OpenStudy (babyslapmafro):

no prob